OpenStudy (anonymous):

What are the vertices of the hyperbola given by the equation ((x-3)^2)/(4) - ((y+7)^2)/(1)=1 ?

OpenStudy (anonymous):

the center you can read off from the equation, it is \((3,-7)\) and we know it looks like this 

OpenStudy (anonymous):

ok crappy picture, but you get the idea. i t looks like that because the denominator of the \(x\) term is larger then the denominator of the \(y\) part, and that also gives you the vertex

OpenStudy (anonymous):

since \(a^2=4\) we know \(a=2\) so you need to go two units right and left of the center namely \((1,-7)\) and \((5,-7)\)

OpenStudy (anonymous):

thank you so much, I completely understand now. would you mind helping me on a different type of question and working me through it? I've been having trouble with it. What are the vertices of the hyperbola given by the equation 16x2 – y2 + 96x + 10y + 103 = 0?

OpenStudy (anonymous):

yeah we have to write it like the other one

OpenStudy (anonymous):

\[16x^2+96x-y^2+10y=-103\] is a start then we have to "complete the square"

OpenStudy (anonymous):

\[16(x^2+6x)-(y^2-10y)=-103\] \[16(x+3)^2-(y-5)^2=-103-5^2+16\times 3^2\] is how you complete the square

OpenStudy (anonymous):

you get \[16(x+3)^2-(y-5)^2=16\] divide by \(16\) and get \[(x+3)^2-\frac{(y-5)^2}{16}=1\] and now we proceed as before

OpenStudy (anonymous):

center is \((-3,5)\) one unit to the left gives \((-4,5)\) and one unit to the right gives \((-2,5)\)

OpenStudy (anonymous):

btw i said something incorrect earlier, you know it looks the way it does because the x part is first and the y part is second, not because of the denominators i was thinking about an ellipse

OpenStudy (anonymous):

\[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] always looks like this 

OpenStudy (anonymous):

do they really always appear like that?